Oakwood Company produces maple bookcases to customer order. The following information is available for the production of theOakwood Company produces maple bookcases to customer order. The following information is available for the production of theOakwood Company produces maple bookcases to customer order. The following information is available for the production of the
Answer #1

Answer to requirement 1 -

Manufacturing cycle time refers to time taken for production of a product right from the start of the production till the time the product is ready to sale. It includes both value added and non value added activities in production of a product. Thus manufacturing cycle time refers to total time required for production of a product in its production cycle.

Thus, Manufacturing cycle time = Process time + Inspection time + Move time + Wait time

Accordingly, in the given case the manufacturing cycle time of Ooakwood company for manufacture of bookcases is as under,

Manufacturing cycle time (days)= 7.0+1.8+4.3+13
  Manufacturing cycle time (days) = 26.1

Answer to requirement 2 -

Manufacturing cycle efficiency refers to a measure that helps to understand the proportion of time spend on value added activity for production of a product. Accordingly,

Manufacturing cycle efficiency = Value added activity time(Only process time is value added activity in production) / Total manufacturing cycle time(from requirement 1)
Manufacturing cycle efficiency = (7.0/26.1)*100
Manufacturing cycle efficiency = 26.82%

Note: Manufacturing cycle efficiency should be always expressed in percentage.

Answer to requirement 3 -

To calculate new manufacturing cycle efficiency by implementing new predictions,
Move time = 3.2 days (4.3 -1.1)
Wait time = 10.5 days (13-2.5)

New Manufacturing cycle time = Process time + Inspection time + Move time + Wait time
New Manufacturing cycle time = 7+1.8+3.2+10.5
New Manufacturing cycle time = 22.5 days

Manufacturing cycle efficiency = 7/22.5(as above)*100
Manufacturing cycle efficiency = 31.11%